An ocean-depth sounding device emits a signal 38kHz in water whose average temperature is 4 degree celcius. The impulse is reflected from the ocean bed and returned to the same location 1.62 seconds after the signal is emmited. Calculate:
(i) the depth of the water
(ii) the wavelength of the signal in water
1
Expert's answer
2020-09-23T09:00:58-0400

An empirical equation for the speed of sound in sea water is the following (see https://en.wikipedia.org/wiki/Speed_of_sound#Water):


"c(T) = a_1 + a_2 T + a_3 T^2 + a_4 T^3"

where "a_1 = 1448.96, a_2 = 4.591, a_3 = -5.304 \times 10^{-2}, a_4 = 2.374 \times 10^{-4}". Thus, for "T = 4\degree C" :


"c(T) = 1448.96 + 4.591\cdot 4 + -5.304 \times 10^{-2}\cdot 4^2 + 2.374 \times 10^{-4}\cdot 4^3 =\\
\approx 1467m/s"

Then, in "t = 1.62s\/2 = 0.81s" the impulse will reach the ocean bed. Thus, it will travell the following distance:


"h = ct = 1467m/s\cdot 0.81 = 1188.27m"

The wavelength is given by the following expression:


"\lambda = \dfrac{c}{\nu}"

where "\nu = 38kHz = 38\times10^3Hz" is the frequency of the wave. Obtain then:

"\lambda = \dfrac{1467m/s}{38\times10^3Hz} = 38.6\times 10^{-3}m"

Answer. (i) "1188.27m", (ii) "38.6\times 10^{-3}m".